# A Simple yet Powerful Instance-Aware Prompting Framework for Training-free Camouflaged Object Segmentation

### v2

@@ Abstract
Camouflaged Object Segmentation (COS) remains highly challenging due to the intrinsic visual similarity between target objects and their surroundings. While training-based COS methods achieve good performance, their performance degrades rapidly with increased annotation sparsity. To circumvent this limitation, recent studies have explored training-free COS methods, leveraging the Segment Anything Model (SAM) by automatically generating visual prompts from a single task-generic prompt (e.g., ”camouflaged animal”) uniformly applied across all test images. However, these methods typically produce only semantic-level visual prompts, causing SAM to output coarse semantic masks and thus failing to handle scenarios with multiple discrete camouflaged instances effectively. To address this critical limitation, we propose a simple yet powerful Instance-Aware Prompting Framework (IAPF), the first training-free COS pipeline that explicitly converts a task-generic prompt into fine-grained instance masks. Specifically, the IAPF comprises three steps: (1) Text Prompt Generator, utilizing task-generic queries to prompt a Multimodal Large Language Model (MLLM) for generating image-specific foreground and background tags; (2) Instance Mask Generator, leveraging Grounding DINO to produce precise instance-level bounding box prompts, alongside the proposed Single-Foreground Multi-Background Prompting strategy to sample region-constrained point prompts within each box, enabling SAM to yield a candidate instance mask; (3) Self-consistency Instance Mask Voting, which selects the final COS prediction by identifying the candidate mask most consistent across multiple candidate instance masks. Extensive evaluations on standard COS benchmarks demonstrate that the proposed IAPF significantly surpasses existing state-of-the-art training-free COS methods.
Codehttps://github.com/ycyinchao/IAPF
Introduction
@@
Camouflaged Object Segmentation (COS) holds significant research value and broad application prospects, including agriculture, medicine, and military reconnaissance. It remains a challenging yet essential task in computer vision, focusing on accurately identifying and segmenting objects visually indistinguishable from their surroundings. Traditional fully-supervised COS methods (Fan et al. 2020; Yin et al. 2024) heavily rely on dense pixel-level annotations, significantly escalating annotation costs and limiting practical applicability. To alleviate annotation burdens, weakly-supervised and unsupervised paradigms (Chen et al. 2024b; Du et al. 2025) have been introduced; however, these training-based methods typically suffer from performance degradation as label sparsity increases.
@@
Recent advancements in promptable segmentation models (e.g., Segment Anything Model, SAM (Kirillov et al. 2023)) have demonstrated promising potential, achieving competitive results across various segmentation tasks with minimal manual visual prompts (i.e., points and boxes). Nevertheless, manual prompts significantly constrain their practical deployment when segmenting specific objects of interest. Consequently, increasing attention has been devoted to developing training-free segmentation methods that automatically generate visual prompts without additional supervision. An active research area within this context is training-free COS methods (Yin et al. 2025; Hu et al. 2024a, b; Tang et al. 2024), which utilizes a single task-generic prompt (e.g., ”camouflaged animal”) indiscriminately applied across all test samples within the COS domain to automatically generate visual prompts required by SAM, thus enabling segmentation of camouflaged objects of interest.
Despite such advancements, existing methods primarily produce semantic-level visual prompts rather than instance-level prompts, directly determining whether SAM outputs semantic or instance masks. According to SAM, explicit bounding boxes are required for instance segmentation. Thus, intuitively, generating instance-level bounding boxes is crucial. Unfortunately, current methods fail to meet this requirement adequately: 1) As illustrated in Fig. 1(a)(1), existing approaches employing Multimodal Large Language Models (MLLM) (Liu et al. 2024) via visual question-answering typically produce a single box covering all objects collectively. 2) Alternatively, methods using intermediate semantic masks (Fig. 1(a)(2)) adopt a MaxIOUBox approach, which identifies all connected components in the mask, computes bounding boxes, and retains the largest one. The above two approaches inevitably result in a single box per image, which typically covers only one or overlapping multiple instances, severely limiting their efficacy in scenarios containing multiple discrete camouflaged instances.
As demonstrated in Fig. 1(a)(3), our method leverages the instance-aware detection capability of Grounding DINO (Liu et al. 2023) to generate multiple instance-level bounding boxes. This enables SAM to perform fine-grained instance segmentation even in complex camouflage scenes. To quantify the accuracy of bounding box prompts generated by the above methods, we evaluate their performance in Fig. 1(b). It is evident that our approach significantly outperforms existing methods in bounding box prompts accuracy on the COD10K dataset (Fan et al. 2020), achieving a mAP of $0.662$ at a stringent IoU threshold of $0.75$, substantially surpassing the $0.009$ and $0.065$ achieved by the above two methods, respectively.
@@
To overcome these limitations, we propose a simple yet powerful Instance-Aware Prompting Framework (IAPF) specifically designed for training-free Camouflaged Object Segmentation. As depicted in Fig. 2, the IAPF initially utilizes an MLLM to automatically generate sets of foreground and background category tags from the input image. Subsequently, leveraging these text prompts, the proposed Instance Mask Generator incorporates Grounding DINO (Liu et al. 2023) to accurately produce multiple instance-aware bounding box prompts, overcoming the critical shortcomings of existing methods. Additionally, within the Instance Mask Generator, we also introduce a region-constrained point prompting strategy named Single-Foreground Multi-Background Prompting (SFMBP), which efficiently leverages multiple background tags to further refine point prompts for instance segmentation. These carefully selected visual prompts (i.e., points and boxes) enable SAM to generate a precise and fine-grained candidate instance mask. Lastly, we introduce a Self-consistency Instance Mask Voting mechanism by evaluating multiple candidate instance masks, ensuring final robust and accurate COS results. Our primary contributions are summarized as follows:
@@
    We present a novel, training-free Instance-Aware Prompting Framework that effectively addresses the limitations of semantic-level visual prompts in existing training-free COS methods, enabling accurate and complete Camouflaged Object Segmentation.
@@
    We propose the Instance Mask Generator, an integral module that combines Grounding DINO and our SFMBP strategy to generate precise instance-level visual prompts, substantially improving the segmentation completeness and accuracy in complex camouflage scenarios.
@@
    Extensive experimental evaluations on multiple COS benchmarks demonstrate that the proposed IAPF significantly outperforms existing training-free COS methods.
@@
Figure 2: Framework of the proposed IAPF, which consists of three steps: (1) Text Prompt Generator: an MLLM turns task-generic prompts and an image into foreground/background tags. (2) Instance Mask Generator (the lower part): Grounding DINO produces NN bbox prompts; the Single-Foreground Multi-Background Prompting (SFMBP) strategy uses CLIP to create foreground and multi-background heatmaps and samples region-constrained point prompts; SAM converts NN visual prompt pairs (i.e., points and boxes) to a candidate instance mask. (3) Self-consistency Instance Mask Voting mechanism: The instance mask with the highest consistency is voted as the final COS prediction.
Related Work
Camouflaged Object Segmentation
Early fully-supervised COS networks (Fan et al. 2020; Hu et al. 2023) achieve high accuracy by learning from dense pixel masks, but these masks are expensive to obtain, and every new domain requires additional annotation and retraining. Weakly-supervised alternatives (Chen et al. 2024a; He et al. 2023b; Li et al. 2025a) alleviate the burden by using sparse cues—-points, scribbles, or boxes—-yet their accuracy drops rapidly as the cues become scarce and they still need task-specific optimisation. Unsupervised pipelines (Zhang and Wu 2023; Shou et al. 2025; Du et al. 2025) remove labels altogether, but at the cost of a clear performance gap. In short, all existing COS families are training-centric, and their effectiveness is tightly coupled to the amount of supervision available, motivating a paradigm that can segment camouflaged objects without any training or annotation.
Segment Anything Model for COS
SAM (Kirillov et al. 2023) produces precise masks only when supplied with explicit visual prompts. Early COS adaptations fine-tune SAM with dense (fully-supervised (Chen et al. 2023; Gao et al. 2025)) or sparse (weakly-supervised (He et al. 2023a; Chen et al. 2024b)) labels, but these variants still require training. To remove this dependency, training-free pipelines automatically generate visual prompts for SAM at test time. GenSAM (Hu et al. 2024a) is the first to do so: it lets an MLLM produce text prompts, feeds them to CLIP to obtain visual prompts for SAM. ProMaC (Hu et al. 2024b) improves GenSAM by exploiting hallucination priors to reduce the inaccuracy of MLLM-generated text prompts. RDVP-MSD (Yin et al. 2025) targets the other source–noisy CLIP heatmaps–by constraining point sampling to a box, yet it still relies on a single semantic-level box prompt. Consequently, all these methods struggle when an image contains multiple camouflaged objects.
Method
Framework Overview
@@
We propose the Instance-Aware Prompting Framework (IAPF), a training-free solution that segments multiple unseen classes of camouflaged objects from a single task-generic prompt. The key novelty is to supply instance-level visual prompts to SAM for predicting instance masks rather than coarse semantic masks. Given a test image $X$ and a task-generic prompt $p_{i}^{g}$ (e.g., ”camouflaged animal”), the IAPF proceeds in three frozen steps. (1) Text Prompt Generator. An MLLM answers a small set of text queries derived from multiple task-generic prompts $p_{i}^{g}$ and returns foreground and background category tags, noted as $tag^{fg}$ and $tag^{bg}$. (2) Instance Mask Generator. Grounding DINO converts each foreground tag into multiple instance-aware bounding boxes, while the proposed Single-Foreground Multi-Background Prompting (SFMBP) strategy samples region-constrained foreground and multi-background points inside every box from CLIP-derived heatmaps; all box-and-points pair forms multiple instance-level visual prompts for SAM, which produces a candidate instance mask. (3) Self-consistency Instance Mask Voting strategy. Among all candidate instance masks, the one closest to their mean is selected as the final COS result. With all models kept frozen, the IAPF delivers instance-aware, zero-shot COS at negligible additional cost.
Text Prompt Generator
The Text Prompt Generator leverages the advanced visual question capabilities of the Multimodal Large Language Models (MLLM) (Liu et al. 2024) to convert a task-generic prompt $p_{i}^{g}$ to text prompts tailored for each input image. Formally, given an input image $X\in\mathbb{R}^{H\times W\times 3}$ representing the visual content, a textual query $Q$ is fed into the MLLM with parameters $\theta$ to generate a textual response sequence $y$. The generated response $y$ explicitly characterizes details relevant to the query. The generation process follows an auto-regressive probabilistic modeling framework, sequentially predicting tokens based on the visual context provided by $X$, the textual query $Q$, and previously generated tokens. Mathematically, the entire response generation can be represented by the following joint probability:
@@
where $y_{t}$ denotes the token generated at time step $t$, $y_{<t}$ represents the sequence of tokens generated before time step $t$, and $T$ is the total length of the generated response.
For simplicity, we adopt the same protocol of GenSAM (Hu et al. 2024a) for extracting the text prompts required by the Instance Mask Generator. Concretely, for each task-generic prompt $p^{g}_{i}$ we construct two textual queries, $Q_{i}^{f}$ and $Q_{i}^{b}$, and submit them to a frozen MLLM:
@@
where $C_{i}$ is the image caption first produced by the MLLM. The two textual queries are instantiated as $Q_{i}^{f}\!:$ “Name of the $p^{g}_{i}$ in one word.” and $Q_{i}^{b}\!:$ “Name of the environment of the $p^{g}_{i}$ in one word.”. The single-word answers returned by the model are recorded as the foreground tag $tag_{i}^{fg}$ and background tag $tag_{i}^{bg}$, respectively.
Instance Mask Generator

### v3

# An Instance-Aware Prompting Framework for Training-free Camouflaged Object Segmentation

@@ Abstract
Training-free Camouflaged Object Segmentation (COS) seeks to segment camouflaged objects without task-specific training, by automatically generating visual prompts to guide the Segment Anything Model (SAM). However, existing pipelines mostly yield semantic-level prompts, which drive SAM to coarse semantic masks and struggle to handle multiple discrete camouflaged instances effectively. To address this critical limitation, we propose an Instance-Aware Prompting Framework (IAPF) tailored for the first training-free COS that upgrades prompt granularity from semantic to instance-level while keeping all components frozen. The centerpiece is an Instance Mask Generator that (i) leverages a detector-agnostic enumerator to produce precise instance-level box prompts for the foreground tag, and (ii) introduces the Single-Foreground Multi-Background Prompting (SFMBP) strategy to sample region-constrained point prompts within each box prompt, enabling SAM to output instance masks. The pipeline is supported by a simple text prompt generator that produces image-specific tags and a self-consistency vote across synonymous task-generic prompts to stabilize inference. Extensive evaluations on three COS benchmarks, two CIS benchmarks, and two downstream datasets demonstrate state-of-the-art performance among training-free methods. Code will be released upon acceptance.
1 Introduction
@@
Camouflaged Object Segmentation (COS) holds significant research value and broad application prospects, including agriculture [38, 37], medicine [11], and military reconnaissance [36], yet remains challenging because targets are visually indistinguishable from their surroundings. Fully-supervised COS methods [10, 39] heavily rely on dense pixel annotations, which are costly and limit practical deployment. Weakly-supervised and unsupervised alternatives [2, 7] mitigate labeling, but still degrade as supervision becomes sparse.
@@
Recent advancements in promptable segmentation models (e.g., Segment Anything Model (SAM) [21]) have demonstrated promising potential, achieving competitive results across various segmentation tasks with minimal manual visual prompts (i.e., points and boxes). Nevertheless, manual prompts significantly constrain their practical deployment when segmenting specific objects of interest. Consequently, increasing attention has been devoted to developing training-free segmentation methods that automatically generate visual prompts without additional supervision. An active research area within this context is training-free COS methods [40, 16, 17, 35], which utilize a single task-generic prompt (e.g., ”camouflaged animal”) indiscriminately applied across all test samples within the COS domain to automatically generate visual prompts required by SAM, thus enabling segmentation of camouflaged objects of interest.
Despite such advancements, existing training-free COS pipelines still generate semantic-level visual prompts rather than instance-level ones, which constrains SAM to output coarse semantic masks and prevents reliable handling of multiple discrete camouflaged objects. According to SAM, explicit bounding boxes are required for instance segmentation; therefore, enumerating instance-level boxes is essential. However, current practices either ask an MLLM to return a single coarse box via VQA (Figure 1(a)(1)) or derive a single box from an intermediate semantic mask through MaxIOUBox (Figure 1(a)(2)), both collapsing multiple instances into one region and failing under multi-object camouflage.
Considering the above issues, we adopt an instance-aware prompting strategy: enumerate all camouflaged instances with multiple boxes and constrain point prompting per box to steer SAM from semantic to instance outputs. The box enumerator is detector-agnostic (instantiated with Grounding DINO [26]) and serves only to provide instance-level box prompts (Figure 1(a)(3)). As shown in Figure 1(b), our box prompts achieve mAP 0.6620.662@ IoU $0.75$ on COD10K, far surpassing semantic-level baselines ($0.009$, $0.065$).
@@
Motivated by the above observations, we propose an Instance-Aware Prompting Framework (IAPF) tailored for training-free COS. As depicted in Figure 2, IAPF first uses an MLLM to generate image-specific sets of foreground and background category tags. Leveraging these text prompts, the proposed Instance Mask Generator instantiates a detector-agnostic enumerator to produce multiple instance-aware box prompts, addressing the core shortcomings of prior work. We further introduce a region-constrained point prompting strategy, Single-Foreground Multi-Background Prompting (SFMBP), which leverages multiple background tags to refine point prompts per instance. These carefully selected instance-level visual prompts (i.e., box-points pairs) enable SAM to produce precise, fine-grained candidate instance masks. Finally, we employ a Self-Consistency Instance Mask Voting mechanism over multiple candidates to yield robust and accurate COS results.
Our contributions are threefold:
@@
    We present IAPF that upgrades prompt granularity from semantic-level to instance-level for training-free COS, enabling accurate and complete segmentation under multi-instance camouflage.
@@
    Instance Mask Generator. We pair a detector-agnostic box enumerator (for proposing multiple instance-level box prompts) with SFMBP (for sampling instance-level point prompts), prompting SAM to accurately and completely produce instance masks in multi-instance camouflage.
@@
    Extensive experiments on three challenging COS benchmarks, two Camouflaged Instance Segmentation (CIS) benchmarks, and two downstream application datasets demonstrate that IAPF achieves SOTA performance among training-free methods.
@@
Figure 2: Framework of the proposed IAPF, which consists of three steps: (1) Text Prompt Generator: an MLLM turns task-generic prompts and an image into foreground/background tags. (2) Instance Mask Generator (bottom): a detector-agnostic enumerator produces NN instance-level box prompts; SFMBP uses CLIP to form foreground/multi-background heatmaps and samples region-constrained per-box points; SAM converts NN box–points pairs into a candidate instance mask. (3) Self-consistency Instance Mask Voting: the candidate whose semantic projection is most consistent across repetitions is selected as the final COS prediction (its instance masks provide CIS).
2 Related Work
2.1 Camouflaged Object Segmentation
Early fully supervised COS models [10, 19] achieved strong accuracy with dense pixel masks, but the annotation cost is prohibitive, and every new domain typically requires re-annotation and retraining. Weakly supervised methods [1, 14, 23] reduce labeling by using sparse cues (points/scribbles/boxes), yet performance degrades as supervision becomes scarce and task-specific optimization remains necessary. Unsupervised pipelines [44, 34, 7] remove labels but exhibit a clear gap to fully supervised performance. Overall, these families are training-centric: their effectiveness scales with supervision and per-task fine-tuning. This motivates a training-free alternative that operates entirely with frozen models and zero task-specific labels, which be pursued by converting text cues into effective visual prompts at inference time.
2.2 Segment Anything Model for COS
SAM [21] delivers accurate masks when supplied with explicit visual prompts (points/boxes). COS adaptations fall into two lines: (i) training-based variants that fine-tune SAM with dense [3, 12] or sparse [13, 2] supervision, and (ii) training-free pipelines that synthesize prompts at inference stage. Representative training-free approaches include: GenSAM [16], which uses an MLLM to produce text tags and derives visual prompts via CLIP heatmaps; ProMaC [17], which exploits hallucination priors to stabilize MLLM-generated tags; UpGen [8], which introduces generative priors to enhance training-free localization and prompt formation; INT [18], which performs negative mining to improve the discriminativeness of task-generic promptable segmentation; and RDVP-MSD [40], which mitigates noisy heatmaps by sampling points within a single region-constrained box. Despite these advances, the existing training-free methods remain at the semantic-level (global or single-box prompts), lacking explicit instance enumeration—struggling when an image contains multiple camouflaged objects.
3 Method
3.1 Framework Overview
@@
We propose the Instance-Aware Prompting Framework (IAPF), a training-free pipeline for COS that upgrades prompt granularity from semantic to instance-level while keeping all components frozen (no task-specific training). Given a test image $X$ and a task-generic prompt $p_{i}^{g}$ (e.g., “camouflaged animal”), IAPF repeats the pipeline $I$ times with synonymous prompts $p_{i}^{g}(i=1,...,I)$ and proceeds in three frozen stages. (1) Text Prompt Generator. An MLLM answers a small set of text queries derived from $p_{i}^{g}$ (and its synonymous variants) and returns image-specific category tags: a set of foreground tags $tag^{fg}$ and a set of background tags $tag^{bg}$. (2) Instance Mask Generator. A detector-agnostic box enumerator converts $tag^{fg}_{i}$ into $N$ instance-level boxes prompts $\mathcal{B}_{i}=\{\mathcal{B}_{i,n}\}_{n=1}^{N}$. Conditioned on $tag^{fg}_{i}$ and $set(tag^{bg})$, the Single-Foreground Multi-Background Prompting (SFMBP) strategy derives region-constrained foreground/background point prompts inside each box using CLIP-based heatmaps, yielding per-box points $\{p^{fg}_{i,n},p^{bg}_{i,n}\}$. All box-points pairs $\langle\mathcal{B}_{i,n},\,p^{fg}_{i,n},\,p^{bg}_{i,n}\rangle(n=1,2,...,N)$ form $N$ instance-level visual prompts for SAM, which produces a candidate instance mask for the image. (3) Self-consistency Instance Mask Voting. Running the above with $I$ times, yields multiple candidate instance masks; the final COS result is chosen as the one closest to the set mean.
3.2 Text Prompt Generator
Text Prompt Generator employs a frozen Multimodal Large Language Model (MLLM) [25] to convert a synonymous task-generic prompt $p_{i}^{g}$ (repetition index $i\in\{1,\dots,I\}$) into image-specific category tags tailored for each input image. Formally, given an image $X\in\mathbb{R}^{H\times W\times 3}$ and a textual query $Q$, the MLLM generates a response sequence $y$ in an auto-regressive manner:
@@ 3.2 Text Prompt Generator
where $y_{t}$ is the token at step $t$, $y_{<t}$ are previously generated tokens, $T$ is the response length, and $\theta$ denotes the (frozen) model parameters.
Following the GenSAM protocol [16], for each $p^{g}_{i}$ we first obtain an image caption $C_{i}$, and then sequentially query the frozen MLLM with two textual queries $Q_{i}^{f}$ and $Q_{i}^{b}$:
@@
where $Q_{i}^{f}:$ “Name of the $p^{g}_{i}$ in one word.” and $Q_{i}^{b}:$ “Name of the environment of the $p^{g}_{i}$ in one word.” The MLLM’s single-word answers are recorded as the foreground tag $tag_{i}^{fg}$ and background tag $tag_{i}^{bg}$, respectively. The pair $(tag_{i}^{fg},\,\mathrm{set}(tag^{bg}))$ constitutes the image-specific category tags consumed by the Instance Mask Generator at repetition $i$.
3.3 Instance Mask Generator
Existing training-free COS pipelines [40, 16, 17, 35] typically operate with a single semantic-level prompt (often one box per image), which constrains SAM [21] to coarse semantic masks and fails on multi-instance camouflage. Instance Mask Generator (IMG) bridges the image-specific category tags from the Text Prompt Generator to instance-level visual prompts under the fully frozen setting. For the $i$-th repetition with tags $(tag^{fg}_{i},\,\mathrm{set}(tag^{bg}))$, IMG performs: (1) Box prompting—a detector-agnostic enumerator converts $tag^{fg}_{i}$ into instance-level box prompts $\mathcal{B}_{i}=\{\mathcal{B}_{i,n}\}_{n=1}^{N}$; (2) Point prompting—conditioned on $tag^{fg}_{i}$ and $\mathrm{set}(tag^{bg})$, SFMBP samples region-constrained foreground/background point prompts inside each $\mathcal{B}_{i,n}$ using CLIP-based heatmaps, yielding per-box point prompts $\{p^{fg}_{i,n},\,p^{bg}_{i,n}\}$; (3) Mask prediction—each box-points pair $\langle B_{i,n},\,p^{fg}_{i,n},\,p^{bg}_{i,n}\rangle$ is fed to SAM to produce an instance mask $m_{i,n}$, and the candidate semantic mask for repetition $i$ is $mask_{i}=\bigvee_{n=1}^{N}m_{i,n}$ for subsequent Self-Consistency Instance Mask Voting. This explicit instance-level guidance—multiple boxes with per-box points—enables robust segmentation of multiple camouflaged objects while keeping all components frozen (no task-specific training).
3.3.1 Box Prompting
Prior training-free COS pipelines [16, 17, 18, 40] often collapse to a single semantic box per image, which prevents SAM [21] from producing true instance masks. We cast box prompting as a detector-agnostic, text-conditioned instance enumeration problem under frozen weights: given the image $X$ and the foreground tag $tag^{fg}_{i}$ from the Text Prompt Generator, a box enumerator converts these into multiple instance-level box prompts:
$\hat{\mathcal{B}}_{i}\;=\;\mathcal{E}_{\text{Detector}}(X,\,tag^{fg}_{i}),$ (4)
where $\mathcal{E}_{\text{Detector}}$ is a plug-and-play open-vocabulary detector (instantiated with Grounding DINO [26] in our implementation). To prevent adjacent camouflaged objects from being merged, we apply a non-maximum suppression (NMS) with a deliberately low IoU threshold,
$\mathcal{B}_{i}\;=\;\mathrm{NMS}_{\text{IoU}=0.2}\!\left(\hat{\mathcal{B}}_{i}\right)\;=\;\{\mathcal{B}_{i,n}\}_{n=1}^{N},$ (5)
where suppresses redundant overlaps while preserving nearby instances. This step upgrades the prompt granularity from semantic to instance by enumerating per-image boxes in a training-free manner, replaces the “single-box” assumption with detector-driven coverage of all camouflaged instances, and yields a measurable intermediate (box-prompt accuracy) to assess prompt quality, as shown in Figure 1(b).
3.3.2 Point Prompting
Motivation. Prior training-free COS methods either draw points from image-level heatmaps [16, 17] or confine sampling to one region-constrained (RC) box per image [40]. Both choices are brittle for multi-instance camouflage: (i) backgrounds are heterogeneous (foliage/sand/rock, etc.), so a single background tag provides weak negatives and frequent confusion; (ii) the single-box setting collapses instances and prevents per-instance score calibration, so global thresholds either miss small objects or overfire around large ones. Semantic vs. instance level: earlier methods yield semantic-level point prompts (global or single-region), whereas we explicitly generate instance-level point prompts per enumerated box to align supervision with each camouflaged instance.
SFMBP (ours). We introduce Single-Foreground Multi-Background Prompting, an instance-aware, multi-background scheme operating per enumerated box $\mathcal{B}_{i,n}$ at repetition $i$, using the foreground tag $tag^{fg}_{i}$ and multiple background tags $\text{set}(tag^{bg})$. Specifically, Spatial CLIP [16] is used to generate a foreground heatmap $H^{f_{i}}$ and multiple background heatmaps $H^{b}=\{H^{b_{1}},H^{b_{2}},\dots,H^{b_{L}}\}$, each $H^{b_{l}}$ corresponds to a specific background tag $tag^{b_{l}}$ from $\text{set}(tag^{bg})$,where $\text{set}(\cdot)$ denotes deduplication over the background tag list before heatmap computation, and $L$ is the length of $\text{set}(tag^{bg})$:
$H^{f_{i}}$ $=\text{Spatial CLIP}(X,tag_{i}^{fg}),$ (6)
$H^{b_{l}}$ $=\text{Spatial CLIP}(X,tag^{b_{l}}),\quad l=1,2,\dots,L$ (7)
and we perform box-wise normalization to calibrate scores locally within each instance. High-confidence, per-box point prompts are then obtained by thresholding (optionally inside a slightly eroded box to avoid boundary ambiguity):
$p_{i,n}^{fg}$ $=\left\{(x,y)\mid\mathrm{Norm}(H^{f_{i}}\mid\mathcal{B}_{i,n})_{(x,y)}\geq 0.9\right\},$ (8)
$p_{i,n}^{\mathrm{bg}}$ $=\bigcup_{l=1}^{L}\left\{(x,y)\mid\mathrm{Norm}(H^{b_{l}}\mid\mathcal{B}_{i,n})_{(x,y)}\geq 0.9\right\},$ (9)
where $\mathrm{Norm}(H\mid\mathcal{B}_{i,n})$ denotes normalization of heatmap $H$ within box $\mathcal{B}_{i,n}$ to $[0,1]$ and the union $\bigcup_{l=1}^{L}$ aggregates per-tag background candidates into a single instance-level background point prompts. SFMBP thus upgrades point prompting along two grounded axes: instance locality (box-wise normalization/selection) and multi-background contrast (aggregating cues from multiple backgrounds), yielding compact yet discriminative instance-level point prompts for SAM.